OpenStudy (anonymous):

The perimeter of kite LMNO is 118 feet. Side MN = 2x + 10 and side NO = 4x - 17. What are the sides lengths of the kite?

OpenStudy (anonymous):

pic

OpenStudy (anonymous):

P=2*MN+2*NO 118=2(2x+10)+2(4x-17)=4x+20+8x-34=12x-14 118=12x-14 12x=132 X=132/12=11 Check: 118=2(2x+10)+2(4x-17) 118=2(22+10)+2(44-17) 118=2(32)+2(27) 118=64+54 True

OpenStudy (anonymous):

I forgot the length of sides: 2 are 32 and 2 are 27

OpenStudy (anonymous):

x = 11 MN = 32 NO = 27